OpenStudy (anonymous):

A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? 112 288 360 432

OpenStudy (anonymous):

\[x = 0.8x + 721x = 0.8x+ 721x - 0.8x = 720.2x = 72x = 72/0.2x = 360 \]

OpenStudy (anonymous):

= 360*

OpenStudy (anonymous):

Do you need an explanation or you got it?